Solve the following system of equations:

 

6x – 3y + 10 = 0                 

2x + y + 9 = 0

Grade:12th pass

3 Answers

Arun
5 years ago
Dear student
 
Multiply 2nd equation by 3 and add it with 1st equation
12 x + 37 = 0
x = -37/12
And 
-74/12 + y + 9 = 0
y = -34/12
Regards
Arun (askIITians forum expert)
Aman Sharma
5 years ago
Take the two equations to be A and B. Now multiply equation B with 3 and add equation A with the new equation to get the value of x. Put the value of x to get the value of y.Answer- x=-37/2,y=-17/6
Aman Sharma
5 years ago
Take the two equations to be A and B. Now multiply equation B with 3 and add equation A with the new equation to get the value of x. Put the value of x to get the value of y.Answer- x=-37/12,y=-17/6. Ignore the previous answer